Write the equation of a line with slope m = 6 and including point (0, 1).
Please help !!

Answer by josgarithmetic(39618)

Slope-Intercept form is y=mx+b and from it, b=y-mx. You just substitute your given point and slope to find b.

b=1-6*0=1, which you could already identify from the given point.

ANSWER: y=6x+1